# Trust-informed Decision-Making Through An Uncertainty-Aware Stacked Neural Networks Framework Case Study in COVID-19 Classification

### v1

@@ VI Results and Discussion
Figure 4: Comparison of Uncertainty-Informed Criteria Across Pre-trained Models Using different UQ Methods at Confidence Threshold 0.1
TABLE IV: Comparative performance of proposed method: Summary of trust-informed criteria across various pre-trained models at the confidence threshold 0.1.
Pretrained UQ Setting CAR CPR TPR FTR RAR MRR TRR FRR
BiT MCD % 37.064 ± 4.997 % 40.731 ± 5.441 % 95.485 ± 0.768 % 0.440 ± 0.148 % 61.201 ± 5.090 % 83.625 ± 1.055 % 14.079 ± 0.842 % 2.297 ± 0.455
BiT Ensemble % 30.663 ± 1.074 % 33.502 ± 1.153 % 95.039 ± 0.844 % 0.178 ± 0.025 % 67.739 ± 1.005 % 85.633 ± 0.483 % 12.249 ± 0.306 % 2.119 ± 0.350
BiT EMCD % 37.865 ± 4.583 % 41.617 ± 5.036 % 95.372 ± 0.872 % 0.490 ± 0.195 % 60.299 ± 4.795 % 83.328 ± 1.079 % 14.214 ± 0.956 % 2.459 ± 0.442
E.NetB0 MCD % 34.319 ± 3.841 % 38.055 ± 4.215 % 94.918 ± 0.922 % 0.476 ± 0.151 % 63.850 ± 3.973 % 82.980 ± 0.855 % 14.681 ± 0.641 % 2.339 ± 0.468
E.NetB0 Ensemble % 25.071 ± 1.021 % 27.607 ± 1.100 % 93.831 ± 1.079 % 0.129 ± 0.025 % 73.284 ± 0.954 % 85.335 ± 0.450 % 12.362 ± 0.212 % 2.303 ± 0.376
E.NetB0 EMCD % 31.305 ± 5.702 % 34.733 ± 6.272 % 94.002 ± 0.862 % 0.416 ± 0.184 % 66.717 ± 5.954 % 83.502 ± 1.130 % 14.272 ± 0.896 % 2.226 ± 0.556
ViT MCD % 46.845 ± 5.484 % 51.126 ± 5.940 % 96.907 ± 0.511 % 0.462 ± 0.165 % 51.665 ± 5.613 % 82.605 ± 1.471 % 15.452 ± 1.255 % 1.943 ± 0.495
ViT Ensemble % 40.995 ± 1.171 % 44.378 ± 1.248 % 97.222 ± 0.487 % 0.173 ± 0.029 % 57.834 ± 1.167 % 85.156 ± 0.434 % 12.885 ± 0.308 % 1.959 ± 0.363
ViT EMCD % 46.672 ± 4.445 % 50.901 ± 4.767 % 96.861 ± 0.595 % 0.467 ± 0.134 % 51.816 ± 4.568 % 82.797 ± 1.055 % 15.228 ± 0.913 % 1.975 ± 0.382
Having established that both the base and meta-models individually exhibit strong performance, it is essential to evaluate the effectiveness of the overall framework when considering the combined outputs of both models. This comprehensive performance assessment is presented in Table IV, which details the proposed metrics discussed in Section III-C.
To start, let’s first examine the results of the ViT model with MCD, providing an interpretation of each metric. For the ViT model using the MCD setting, the CAR is 46.85%. This indicates that 46.85% of the total outcomes were both correct and made with high confidence (i.e., trustworthy). This metric suggests that a significant portion of the model’s predictions can be trusted without further verification.
The CPR is about 51.13%, suggesting that 51.16% of correctly predicted by the base model flagged as trustworthy by meta-model. This suggests that while almost half of the correct predictions made by the base model were identified as trustworthy by the meta-model, the other half were not flagged as trustworthy. This could be due to either the miscalibration of the base model (i.e., correct predictions with a predictive entropy greater than the confidence threshold) or errors from the meta-model.
The TPR is about 96.90%, implying among the instances the model flagged as trustworthy, 96.70% were indeed trustworthy (or in other words correct and confidently made). In other words, the model’s trustworthiness flagging mechanism is highly reliable, as nearly all of the predictions marked as trustworthy were both accurate and confident.
The FTR 0.462% indicates that 0.46% of the total outcomes erroneously flagged incorrect predictions as trustworthy. While this error rate is quite low, it underscores the potential for further reducing the chances of misplaced trust in incorrect outcomes.
The RAR is around 51.67%, meaning that over half of the model’s predictions were flagged as untrustworthy (i.e., either uncertainly correct, uncertainly incorrect, or confidently incorrect). This high percentage leading to a significant review burden.
The MRR is 82.60%, indicating that 82.60% of the instances flagged as untrustworthy for review (in other words, 82.60% of RAR) are due to the base model’s miscalibration, where correct predictions with low confidence (or high uncertainty) are flagged as untrustworthy and require further review. This high percentage indicates that the base model struggles to be confident in its correct predictions.
The TRR is relatively low at 15.45%, showing that 15.45% of RAR were correctly flagged as untrustworthy due to incorrect predictions. The FRR is low at around 1.94%, indicating that a small percentage of RAR were mistakenly flagged as untrustworthy due to meta-model errors (flag confidently correct prediction as untrustworthy by meta-model error). This is a positive outcome, as it suggests the model rarely creates unnecessary reviews by incorrectly doubting its correct predictions.
Across all UQ settings, it is apparent that there exists a notable relationship between the CAR and RAR. As CAR increases (from 25.07% with EfficientNetB0 Ensemble to 46.85% with ViT MCD), there is a concomitant decrease in RAR (from 73.28% to 51.67%). This trend aligns with the intuitive understanding that enhancing model performance inherently reduces the need for additional review, thereby diminishing the review burden.
The MCD setting, with a higher AUC score, excels in distinguishing correct predictions with confidence. This is reflected in its relatively higher CAR and CPR values. On the other hand, the Ensemble setting, despite having almost similar AUC scores to EMCD, excels with the lowest FTR. This means that the Ensemble setting by aggregating predictions from multiple models enhances the F1 score, thereby reducing FTR. However, Ensemble’s lower CAR and CPR, despite its high F1 score, suggest that while it can identify correct predictions its confidence in these predictions is not as well-calibrated due to the variability in predictions introduced by the ensemble approach. This variability in predictions observed in the Ensemble setting is also reflected in the higher RAR and MRR. Conversely, MCD, by incorporating dropout during inference, effectively captures uncertainty and enhances the model’s ability to discern when it should be confident in its predictions. However, the slightly higher FTR compared to Ensemble suggests that MCD may occasionally place trust in incorrect predictions. EMCD, by introducing stochasticity during the inference phase of the ensemble, provides improved uncertainty estimations. This approach achieves an AUC comparable to that of Ensemble while attaining higher CAR and CPR, along with lower RAR and MRR, demonstrating performance closely rivaling that of MCD. However, it is important to note that while EMCD and MCD exhibit similar performance, EMCD is considerably more computationally demanding due to the necessity of maintaining an ensemble of models (30 in this study), as opposed to the single fine-tuned model employed in MCD. These findings suggest that the MCD approach may offer a more reliable and stable solution, striking an optimal balance between performance efficiency and computational complexity.
Across the evaluated pre-trained models, ViT consistently demonstrates superior performance compared to BiT and EfficientNetB0, particularly in its ability to generate confident and accurate predictions. Generally, EfficientNetB0 lags behind ViT and BiT, showing lower CAR and higher FTR values. This suggests that while EfficientNetB0 can achieve reasonable classification results, it may not capture the nuances of the data as effectively as the other models. These findings underscore the critical role of the pre-trained model in effectively capturing prediction uncertainty and facilitating trust-informed decision-making.
VI-B Ablation study part One: Influence of Variable Confidence Thresholds
To further investigate the impact of the confidence threshold ($\tau$) on the performance of the meta-model, an ablation analysis was conducted. This analysis involved systematically varying $\tau$ from 0.05 to 0.4 and observing the changes in the model’s ability to distinguish between trustworthy and untrustworthy predictions, illustrated by Figure 5 and Table Supplementary S.V. The error-based performance metrics, such as F1 score and AUC, of the meta-model at each confidence threshold are presented in Supplementary Table S.IV. While all meta-models across various confidence thresholds demonstrate strong performance, the focus of this section is to examine the effect of the confidence threshold on trust-informed decision-making.
It is worth recalling that a PE value closer to 1 indicates lower confidence (higher uncertainty), while a PE value closer to 0 reflects higher confidence (lower uncertainty). Accordingly, as the $\tau$ increases, the model becomes less conservative—meaning it is more permissive about what it labels as trustworthy. When the threshold is low (close to zero), the model is stricter, only labeling correct predictions as trustworthy if they have very low entropy, thus demonstrating high confidence. Thus a clear upward trend in the Mean CAR is observed across all pre-trained models and UQ settings. This trend indicates that as the threshold increases, the model becomes more willing to label a greater number of correct predictions as trustworthy, and similarly, CPR also shows an upward trajectory with increasing thresholds.
As the confidence threshold increases, TPR initially rises, reaches a peak, and then either stabilizes or decreases. At lower thresholds (e.g., 0.05), the model is more conservative, labeling only predictions with very low uncertainty as trustworthy, which limits the number of predictions flagged as trustworthy. As the confidence threshold increases, the model becomes less conservative, allowing more predictions to be flagged as trustworthy, thereby initially improving TPR as more correct predictions are included. However, beyond a certain point, the model’s selectivity decreases, leading to a potential increase in incorrect predictions being erroneously flagged as trustworthy as mirrored by the upward trend in FTR. This suggests that while increasing the confidence threshold can initially improve the model’s trustworthiness in its correct predictions, it may also lead to a higher rate of false trust in incorrect predictions.
As the $\tau$ increases, Mean RAR exhibit a clear downward trend across all models and UQ settings. The decline in RAR as $\tau$ increases indicates that fewer predictions are being flagged as uncertain and sent for review. This is likely because as the model becomes more lenient (by allowing predictions with slightly higher entropy to be considered trustworthy), it flags fewer predictions as requiring review, reducing the review burden. Also, the inverse relationship between CAR and RAR is intuitive: as the model becomes better at identifying trustworthy predictions (higher CAR), the need for further human review diminishes (lower RAR).
Simultaneously, the decrease in MRR with an increasing confidence threshold indicates that the proportion of review cases attributed to base model miscalibration is also reducing. This suggests that as the confidence threshold rises, the instances where the base model is correct but uncertain (leading to unnecessary review) become less frequent. This reduction aligns with the decrease in RAR, reinforcing the idea that a higher confidence threshold not only reduces the total number of reviews but also leads to fewer correctly predicted instances being flagged for review.
As the $\tau$ increases, FRR and TRR exhibit an upward trend. At first glance, this may seem counterintuitive, particularly given the simultaneous decrease in the RAR, which might lead to the expectation that FRR and TRR would also decrease. The explanation lies in the interaction between the changing threshold and the total uncertain predictions (TUP). As $\tau$ increases, the model becomes more permissive, leading to fewer predictions being flagged as untrustworthy, which decreases the TUP (Eq. 15). The correctness of a prediction, as determined by comparing the predicted labels to actual labels, remains independent of the confidence threshold. Therefore, the absolute number of incorrect predictions does not change. The increase in TRR as $\tau$ increases can be attributed to the decrease in TUP. As fewer predictions are labeled as untrustworthy, the relative proportion of incorrect predictions (which remain constant in absolute terms) becomes larger. Similarly, the increase in FRR with rising $\tau$ is due to the shrinking pool of TUP, combined with the possibility of increasing meta-model errors, as indicated by the upward trend in FTR.
((a)) CAR
((b)) CPR
((c)) TPR
((d)) FTR
((e)) RAR
((f)) MRR
((g)) TRR
((h)) FRR
@@ VI-B Ablation study part One: Influence of Variable Confidence Thresholds
VI-C Ablation Study part Two: Evaluating the Integration of PE and Image Embeddings
In this ablation study, the impact of incorporating PE as an input to the meta-model was examined. The meta-model was trained under two conditions: one in which PE was included alongside the feature representations from the pre-trained model, and another in which only the feature representations were used as input. This approach allows for an assessment of the significance of PE in enhancing the meta-model’s performance.
Table V presents the error-based performance results of the meta-model excluding PE at $\tau$=0.1. Supplementary Table S.IV provides a summary of the error-based performance results, comparing models with and without PE across various $\tau$ values. Figure 6 illustrates the performance comparison of the meta-model at $\tau$=0.1, while Supplementary Figures S.1 to S.4 display the error-based performance comparisons at $\tau$ values of 0.05, 0.2, 0.3, and 0.4, respectively.
((a)) F1
((b)) AUC
Figure 6: Comparative Performance of Meta-Models With and Without PE as Input at Confidence Threshold 0.1.
@@ VI-C Ablation Study part Two: Evaluating the Integration of PE and Image Embeddings
ViT Ensemble %92.5409 ± 0.4308 %98.2399 ± 0.1796
ViT EMCD %90.9633 ± 1.0366 %97.3828 ± 0.5639
The comparison between meta-models that include PE as input and those that exclude it reveals a consistent trend across both F1 scores and AUC values. Meta-models that incorporate PE as part of their input tend to achieve higher F1 scores, as well as AUC, across all pre-trained models and UQ settings. This suggests that incorporating PE enables the meta-model to more effectively differentiate between trustworthy and untrustworthy predictions. The additional information provided by PE, which quantifies the uncertainty of predictions, allows the model to make more informed decisions. The underlying reason for these improvements is that PE provides the meta-model with crucial contextual information about the certainty of predictions. Without PE, the meta-model relies solely on feature representations, which may not fully capture the uncertainty associated with each prediction. By incorporating PE, the model can adjust its predictions based on its confidence, leading to more accurate and reliable performance. This is particularly valuable in settings where the cost of incorrect predictions is high, and it is essential to identify instances where the model is uncertain.
@@
Improvement in the meta-model’s performance also enhances trust-informed decision-making. Table VI presents the results of the proposed trust-informed metrics excluding PE at $\tau$=0.1. Supplementary Table S.V provides a summary of these results, comparing models with and without PE across various $\tau$ values. Figure 7 illustrates the performance comparison of the proposed metrics at $\tau$=0.1, while Supplementary Figures S.5 to S.8 display the performance comparisons at $\tau$ values of 0.05, 0.2, 0.3, and 0.4, respectively.
The enhanced performance of the meta-model, as demonstrated in the presented results, translates into improved trust-informed decision-making. The improvement is reflected in the increase in CAR, indicating that the model is more accurate when making confident predictions. Similarly, the improvement in CPR with the inclusion of PE indicates that when the model is correct in its predictions, it is more likely to be confident. This is a crucial factor in trust-informed decision-making, as it implies that the meta-model is not only making accurate predictions but is also more reliable when it asserts trust. The enhanced performance of the meta-model is more significant in TPR suggesting that incorporating PE allows the meta-models to better discern between correct predictions that are genuinely confident and those that are not. The increase in CPR and TPR across different settings further supports the notion that PE helps the model to align its confidence with its accuracy, making the overall decision-making process more trustworthy.
Including PE appears to lower the FTR across various pre-trained models and UQ settings. This result is expected because FTR reflects instances where both the base model and the meta-model make errors. By enhancing the meta-model’s performance with the inclusion of PE (as reflected in Figure 6, the likelihood of errors by the meta-model decreases. As a result, the overall incidence of compounded errors—where both models fail simultaneously—decreases.
The RAR, which measures the proportion of predictions flagged as untrustworthy, appears relatively unaffected by the inclusion or exclusion of PE across different settings. This stability in RAR can be attributed to the balancing effect within the total uncertain predictions (TUP) metric. On one hand, the inclusion of PE enhances the meta-model’s ability to correctly identify trustworthy (TT) versus untrustworthy (TU) predictions. This improvement should theoretically decrease the number of incorrect predictions flagged as untrustworthy (FU), which would reduce the TUP and, by extension, the RAR. On the other hand, the same improvement might also enhance the meta-model’s ability to correctly identify uncertainly correct predictions (TPTU, TNTU), which could increase these components of the TUP. As a result, the overall TUP might not change significantly.
The exclusion of PE appears to result in a slight decrease in MRR, particularly noticeable in ensemble settings. This decrease, however, does not necessarily indicate an actual improvement. Rather, it may reflect the meta-model’s reduced ability to accurately discern between trustworthy and untrustworthy predictions. MRR is calculated based on how well the model identifies instances where the base model is either confidently incorrect (FPTT, FNTT) or uncertainly correct (TPTU, TNTU). When PE is excluded, the meta-model’s performance decreases, making it less effective in correctly identifying these instances. As a result, the MRR metric may decrease, not because the base model is better calibrated, but because the meta-model is less capable of capturing true instances of miscalibration. Therefore, the slight decrease in MRR when excluding PE suggests a reduction in the meta-model’s effectiveness in highlighting the base model’s calibration issues, rather than an actual improvement in calibration.
The inclusion of PE results does not significantly impact the TRR. TRR measures the proportion of instances where the model makes an incorrect prediction and correctly flags this prediction as untrustworthy. The observed pattern where including PE does not significantly alter TRR suggests that PE might not be adding much value in terms of enhancing the meta-model’s ability to recognize these erroneous predictions. This could be because the TRR is heavily dependent on the base model’s performance since it is based on the incorrect predictions made by the base model. If the base model’s error patterns are consistent and not highly correlated with PE, the meta-model may find limited use for PE in distinguishing trustworthy from untrustworthy errors. This suggests that while PE can be valuable in other contexts, such as improving other trust-related metrics like TPR and FTR, its role in enhancing TRR specifically is limited in this setup.
On the other hand, the inclusion of PE in the meta-model shows a distinct impact on the FRR. FRR is formulated as the proportion of correct predictions that are incorrectly flagged as untrustworthy due to the meta-model’s error. It is evident that including PE leads to a noticeable decrease in FRR across most pre-trained models and UQ settings. The improvement in the meta-model’s performance ensures that fewer correct predictions are mistakenly flagged for review. In conclusion, the inclusion of PE significantly enhances the meta-model’s ability to discern between trustworthy and untrustworthy predictions, leading to a more reliable and efficient trust-informed decision-making process.
TABLE VI: Summary of trust-informed criteria across various pre-trained models, Excluding PE as Input at the confidence threshold 0.1.
Pretrained UQ Setting CAR CPR TPR FTR RAR MRR TRR FRR
BiT MCD % 35.361 ± 5.106 % 38.868 ± 5.567 % 89.421 ± 1.189 % 0.648 ± 0.292 % 60.415 ± 5.986 % 80.351 ± 2.156 % 13.961 ± 0.833 % 5.688 ± 1.494
BiT Ensemble % 28.736 ± 1.216 % 31.406 ± 1.315 % 90.052 ± 1.560 % 0.232 ± 0.049 % 68.072 ± 1.639 % 83.458 ± 0.660 % 12.153 ± 0.258 % 4.389 ± 0.645
BiT EMCD % 34.646 ± 3.880 % 38.111 ± 4.262 % 89.494 ± 1.259 % 0.597 ± 0.209 % 61.264 ± 4.520 % 79.956 ± 2.052 % 13.918 ± 0.745 % 6.126 ± 1.589
E.NetB0 MCD % 31.121 ± 5.406 % 34.487 ± 5.981 % 87.898 ± 1.831 % 0.661 ± 0.335 % 64.552 ± 6.378 % 80.051 ± 2.315 % 14.181 ± 0.884 % 5.768 ± 1.616
E.NetB0 Ensemble % 24.154 ± 1.270 % 26.592 ± 1.391 % 89.103 ± 1.792 % 0.194 ± 0.044 % 72.870 ± 1.725 % 83.345 ± 0.584 % 12.323 ± 0.375 % 4.331 ± 0.772
E.NetB0 EMCD % 30.786 ± 4.908 % 34.151 ± 5.407 % 87.659 ± 2.128 % 0.636 ± 0.269 % 64.824 ± 5.882 % 79.386 ± 2.257 % 14.309 ± 0.827 % 6.305 ± 1.662
ViT MCD % 45.522 ± 5.137 % 49.690 ± 5.573 % 90.339 ± 1.324 % 0.781 ± 0.266 % 49.571 ± 5.952 % 76.462 ± 2.582 % 15.491 ± 1.222 % 8.047 ± 1.693
ViT Ensemble % 38.329 ± 1.072 % 41.489 ± 1.143 % 90.771 ± 1.096 % 0.272 ± 0.053 % 57.765 ± 1.378 % 81.480 ± 0.884 % 12.720 ± 0.317 % 5.800 ± 0.948
ViT EMCD % 43.850 ± 4.403 % 47.841 ± 4.764 % 89.831 ± 1.294 % 0.704 ± 0.248 % 51.146 ± 5.269 % 77.165 ± 2.199 % 15.042 ± 1.029 % 7.793 ± 1.543
((a)) CAR
((b)) CPR
((c)) TPR
((d)) FTR
((e)) RAR
((f)) MRR
((g)) TRR
((h)) FRR
@@
Figure 7: Comparison of Uncertainty-Informed Criteria Across Pre-trained Models at a Confidence Threshold of 0.1, With and Without PE as Input.

### v2

@@ VI Results and Discussion
Figure 3: Comparison of Uncertainty-Informed Criteria Across Pre-trained Models at a Confidence Threshold of 0.1, U-SNN and Tradiditonal Threshold-based method.
To evaluate the comparative performance of the U-SNN and traditional threshold-based approaches, trends in key metrics were analyzed as the confidence threshold increased from 0.05 to 0.4. Figure 4 illustrates the trends in evaluation metrics for both the proposed U-SNN and the traditional threshold-based approaches across three pre-trained models (BiT, EfficientNetB0, and ViT), with confidence thresholds varying from 0.05 to 0.4. Both methods exhibited a consistent increase in CR as the threshold increased. This is expected, as higher thresholds result in a greater number of predictions being classified as confident. While CR trends were similar, subsequent metrics revealed significant differences in prediction correctness and referral efficiency.
At lower thresholds (0.05 to 0.2), U-SNN consistently outperformed the threshold-based method in minimizing confidently incorrect predictions. For example, BiT’s FCR in U-SNN rose from 0.21% at 0.05 to 1.12% at 0.2, whereas the threshold-based approach increased from 0.74% to 1.38% over the same range. However, at the highest threshold of 0.4, this pattern reversed. The threshold-based approach achieved a lower FCR of 1.99%, compared to U-SNN’s 3.31%. Similar reversals were observed in EfficientNetB0 and ViT.
A similar trend was observed in CE, which also favored U-SNN at lower thresholds but reversed at higher thresholds. For BiT, U-SNN increased CE from 0.69% at 0.05 to 2.09% at 0.2, consistently maintaining lower values than the threshold-based method. However, at 0.4, the traditional approach yielded a lower CE of 2.59%, compared to U-SNN’s 4.35%.
FCR and CE results suggest that as the threshold becomes more lenient, the U-SNN’s performance degrades. This may be due to the meta-model’s diminishing ability to differentiate trustworthy predictions when a large proportion of predictions are automatically deemed confident. Therefore, the U-SNN framework is particularly advantageous in moderate-threshold scenarios, where balancing certainty with correctness is critical. However, when the system operates under very high thresholds, where almost all predictions are labeled as confident, the simpler threshold-based approach may provide tighter control over confidently incorrect predictions.
Although the threshold-based approach exhibited lower FCR and CE at the highest evaluated threshold (0.4), this outcome is less critical when considering practical applications. In uncertainty quantification, thresholds above 0.5 are generally discouraged, as they increase the risk of accepting low-confidence predictions as confident, particularly in high-stakes contexts such as medical diagnosis. Lower thresholds (e.g., 0.05 to 0.2) are scientifically favored as they encourage more conservative and reliable uncertainty estimation. Importantly, U-SNN demonstrated superior performance across these practically relevant lower thresholds, substantially reducing confidently incorrect predictions and redundant referrals. Thus, while U-SNN’s performance slightly degrades at higher thresholds, this scenario is less operationally relevant, and the framework remains superior in the thresholds that matter most for trustworthy decision-making.
In terms of UR and RR, both methods exhibited a decreasing UR trend with rising thresholds, indicating that fewer predictions were flagged for review. This consistent trend suggests that both methods became less conservative in uncertainty classification as the threshold increased. The most notable advantage of U-SNN was observed in the RR metric. Across all thresholds, U-SNN consistently achieved substantially lower RR values. This highlights that U-SNN was consistently more efficient in minimizing unnecessary referrals, ensuring that human reviewers were focused on genuinely uncertain cases. The threshold-based method, however, remained inefficient, flagging a high number of already-correct predictions for unnecessary review.
((a)) CR
((b)) FCR
((c)) CE
((d)) UR
((e)) $R$
@@
VI-A Ablation Study
In this ablation study, the impact of incorporating PE as an input to the meta-model was examined. The meta-model was trained under two conditions: one in which PE was included alongside the feature representations from the pre-trained model, and another in which only the feature representations were used as input. The goal is to determine whether the inclusion of PE contributes to more reliable and efficient uncertainty-aware decision-making within the U-SNN framework.
Table V presents the error-based performance results of the meta-model excluding PE at confidence threshold 0.1. Supplementary Table S.2 provides a summary of the error-based performance results, comparing models with and without PE across various confidence thresholds values.
The comparison between meta-models that include PE as input and those that exclude it reveals a consistent trend across both F1 scores and AUC values. Meta-models that incorporate PE as part of their input tend to achieve higher F1 scores, as well as AUC, across all pre-trained models. The underlying reason for these improvements is that PE provides the meta-model with crucial contextual information about the certainty of predictions. Without PE, the meta-model relies solely on feature representations, which may not fully capture the uncertainty associated with each prediction. By incorporating PE, the model can adjust its predictions based on its confidence, leading to more accurate and reliable performance.
@@ VI-A Ablation Study
Improvement in the meta-model’s performance also enhances trust-informed decision-making. Table VI presents the results of the proposed metrics excluding PE.
To start, let’s first examine the results at confidence threshold 0.1. Figure 5 illustrates the performance comparison of the U-SNN at confidence threshold 0.1. At the confidence threshold of 0.1, including PE in the meta-model consistently resulted in superior performance across several key uncertainty metrics when compared to excluding PE. Excluding PE led to a slight increase in CR across all models. Although a higher CR suggests a larger proportion of predictions being labeled as trustworthy, this increase was marginal. The exclusion of PE resulted in a notable deterioration in FCR. For instance, FCR in the ViT model rose from 0.4623 (with PE) to 0.7815 (without PE), marking a 69.1% increase. Similarly, the BiT model exhibited a 47.4% increase (from 0.4395 to 0.6478). This sharp increase indicates that excluding PE significantly heightens the risk of confidently incorrect predictions, undermining trust in the model’s certainty assessments.
CE also showed a substantial increase when PE was excluded. For example, at confidence threshold 0.1, CE for the ViT model increased by 62.2% (from 0.9315 to 1.5108), while EfficientNetB0 exhibited a 36.5% increase. These results underscore PE’s critical role in helping the meta-model accurately discern between confidently correct and incorrect predictions.
Excluding PE led to a minor reduction in UR across all models. While this reduction suggests fewer referrals, it also implies that some uncertain and potentially incorrect predictions might be wrongly accepted as confident, contributing to the previously noted increase in FCR and CE.
To further understand the broader impact of PE, the analysis was extended across all evaluated confidence thresholds.
Across all thresholds, excluding PE consistently resulted in higher CR values. For example, at 0.4, CR for the ViT model increased from 77.28% (with PE) to 81.41% (without PE). Similar trends were observed for BiT and EfficientNetB0. However, while higher CR suggests greater assertiveness in labeling predictions as trustworthy, it does not inherently indicate reliability, especially when juxtaposed with elevated FCR and CE values.
FCR increased notably when PE was excluded, particularly at lower confidence thresholds. For instance, at 0.05, FCR in ViT increased from 0.2212 (with PE) to 0.4078 (without PE), indicating an 84.4% increase. As thresholds rose, the gap narrowed, but models excluding PE still demonstrated consistently higher FCR values, signaling a persistent vulnerability to confidently incorrect predictions.
The CE metric mirrored FCR trends, with significant increases observed upon PE exclusion. For instance, CE in EfficientNetB0 at 0.2 rose by 30.7%, from 2.3804 (with PE) to 2.9546 (without PE). The highest differences were again observed at lower thresholds, reaffirming the critical role of PE in reducing confidently incorrect predictions in these ranges. UR generally decreased when PE was excluded, particularly at higher thresholds. For instance, UR in the ViT model dropped from 22.72% to 18.59% at the 0.4 threshold. Although a lower UR could suggest fewer unnecessary referrals, this coincided with higher confidently incorrect predictions, raising concerns about the reliability of certainty assessments without PE. RR remained largely stable across all thresholds and models, regardless of PE inclusion. Minor fluctuations were observed, such as a 1.2% increase in RR for the BiT model at 0.4, but these differences were not substantial.
TABLE VI: Performance Summary of U-SNN including and excluding PE as input of meta-model
Pretrained Confidence Threshold U-SNN Input CR FCR CE UR $R$
BiT 0.05 Include PE 29.4514 ± 4.4865 0.2081 ± 0.0749 0.6902 ± 0.1652 70.5486 ± 4.4865 87.4771 ± 0.6375
Exclude PE 29.5792 ± 5.3093 0.3093 ± 0.1735 0.9804 ± 0.3992 70.4208 ± 5.3093 87.5615 ± 0.6489
E.NetB0 Include PE 27.4149 ± 3.7112 0.2182 ± 0.0897 0.7724 ± 0.2322 72.5851 ± 3.7112 86.7379 ± 0.5236
Exclude PE 26.4713 ± 5.6186 0.3159 ± 0.1889 1.1026 ± 0.4884 73.5287 ± 5.6186 87.0982 ± 0.7385
ViT Include PE 39.3139 ± 5.8126 0.2212 ± 0.0981 0.5408 ± 0.1706 60.6861 ± 5.8126 86.4584 ± 1.0562
Exclude PE 40.988 ± 5.8228 0.4078 ± 0.1494 0.964 ± 0.2512 59.012 ± 5.8228 86.3697 ± 1.0404
BiT 0.1 Include PE 38.7992 ± 5.0042 0.4395 ± 0.1458 1.107 ± 0.2486 61.2008 ± 5.0042 85.9213 ± 0.828
Exclude PE 39.5851 ± 5.8851 0.6478 ± 0.2872 1.5713 ± 0.4683 60.4149 ± 5.8851 86.0389 ± 0.8188
E.NetB0 Include PE 36.1502 ± 3.9059 0.4755 ± 0.1488 1.2907 ± 0.2857 63.8498 ± 3.9059 85.3193 ± 0.6299
Exclude PE 35.4484 ± 6.2704 0.6607 ± 0.3294 1.7619 ± 0.6362 64.5516 ± 6.2704 85.8188 ± 0.8694
ViT Include PE 48.3345 ± 5.5188 0.4623 ± 0.1625 0.9315 ± 0.2372 51.6655 ± 5.5188 84.5481 ± 1.2343
Exclude PE 50.4294 ± 5.852 0.7815 ± 0.2616 1.5108 ± 0.3701 49.5706 ± 5.852 84.5088 ± 1.2015
BiT 0.2 Include PE 52.9509 ± 5.2666 1.1213 ± 0.2945 2.0862 ± 0.3533 47.0491 ± 5.2666 83.098 ± 1.1284
Exclude PE 54.2792 ± 5.4427 1.4892 ± 0.4291 2.6953 ± 0.5237 45.7208 ± 5.4427 83.3864 ± 0.9541
E.NetB0 Include PE 50.0759 ± 4.1463 1.2038 ± 0.2583 2.3804 ± 0.339 49.9241 ± 4.1463 82.6654 ± 0.8199
Exclude PE 49.535 ± 6.5867 1.5109 ± 0.5621 2.9546 ± 0.775 50.465 ± 6.5867 83.5164 ± 1.1192
ViT Include PE 60.0983 ± 4.7476 1.0308 ± 0.2663 1.692 ± 0.3233 39.9017 ± 4.7476 81.4207 ± 1.411
Exclude PE 63.2945 ± 5.2023 1.6386 ± 0.4265 2.5508 ± 0.493 36.7055 ± 5.2023 81.4236 ± 1.3043
BiT 0.4 Include PE 75.6858 ± 3.5126 3.3056 ± 0.4536 4.3513 ± 0.4165 24.3142 ± 3.5126 76.3087 ± 1.4808
Exclude PE 78.0664 ± 4.0813 4.1313 ± 0.6528 5.2644 ± 0.5856 21.9336 ± 4.0813 77.4471 ± 1.5522
E.NetB0 Include PE 72.9876 ± 3.8683 3.4535 ± 0.4571 4.7129 ± 0.3947 27.0124 ± 3.8683 76.2099 ± 1.5066
Exclude PE 74.436 ± 5.4932 4.0678 ± 0.9003 5.4068 ± 0.8489 25.564 ± 5.4932 77.4412 ± 1.7935
ViT Include PE 77.2759 ± 3.3395 2.669 ± 0.4377 3.4373 ± 0.4289 22.7241 ± 3.3395 74.6181 ± 1.7885
Exclude PE 81.4089 ± 3.8916 3.7859 ± 0.6173 4.6256 ± 0.557 18.5911 ± 3.8916 74.9084 ± 1.7941
((a)) CR
((b)) FCR
((c)) CE
((d)) UR
((e)) $R$
@@
Figure 5: Comparative Performance of U-SNN With and Without PE as Input at Confidence Threshold 0.1.